OpenStudy (anonymous):

(Complex Numbers) Multiply or divide as indicated: √-7 * √-15

OpenStudy (jdoe0001):

\(\bf \sqrt{-7}\cdot \sqrt{-15}\implies \sqrt{-7\cdot -15}\implies \quad ?\)

OpenStudy (agent0smith):

@jdoe0001 you can't multiply complex numbers that way, you first have to pull out the i \[\large \sqrt{-7}\cdot \sqrt{-15} = i \sqrt 7 *i \sqrt 15\] \[\Large i*i \sqrt 7 \sqrt {15} =\] \[\Large i^2 \sqrt {7*15}\]

OpenStudy (anonymous):

In the answer key, it ends up being: -√105 so does the i^2 become the negative that is before the square root symbol?

OpenStudy (anonymous):

yeah \[ i=\sqrt{-1}\] \[i^2 = -1\]

OpenStudy (anonymous):

How is it a negative though whenever: -1*-1=1

OpenStudy (anonymous):

since \[\sqrt{-1}≠-1 \quad ; \quad \sqrt{-1} = i\] so \[ i^2 = i*i=\sqrt{-1}\sqrt{-1}=(-1)^{1/2}\cdot (-1)^{1/2} =(-1)^1 = -1\]

OpenStudy (jdoe0001):

hmm... ohhh

OpenStudy (jdoe0001):

\(\bf i^2\implies \sqrt{-1}\cdot \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies -1\)

OpenStudy (jdoe0001):

\(\bf \sqrt{-7}\cdot \sqrt{-15}\implies \sqrt{7}\ i\cdot \sqrt{15}\ i\implies \sqrt{105}\ i^2\implies -\sqrt{105}\)

OpenStudy (anonymous):

does that make sense @DemiRae ?

OpenStudy (anonymous):

Yeah that makes sense, I wasn't thinking of them in their square root form. Thank you for helping me with that!